OpenStudy (anonymous):

Find the number of all combinations of 8 objects a, b, c, d, e, f, g, and h taken 7 at a time.

OpenStudy (anonymous):

Are you familiar with the nCr notation?

OpenStudy (anonymous):

noooo we are supposed to use permutations i believe

OpenStudy (anonymous):

No, this is not a permutation problem.

OpenStudy (anonymous):

This is a combination problem.

OpenStudy (anonymous):

ohhhhh yes you ar correct. can you help me with it?

OpenStudy (anonymous):

Are you familiar with nCr and what it means?

OpenStudy (anonymous):

noo i am not

OpenStudy (anonymous):

Ok. Well it is a notation to solve this exact problem.

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

nCr is an expression which tells you how many different combinations of r different things you can make from a pool of n things.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so in this case how big is our pool?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

Correct. And how many things will we be taking from the pool to make our combinations?

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

Correct. So the answer to this problem is 8C7

OpenStudy (anonymous):

But that doesn't help much cause they want a real number.

OpenStudy (anonymous):

So now, let me give you the formula for nCr, and you can use it to find 8C7

OpenStudy (anonymous):

okay! great!

OpenStudy (anonymous):

\[nCr \equiv \frac{n!}{r!(n-r)!} \]

OpenStudy (anonymous):

So can you use that to figure out what 8C7 is?

OpenStudy (anonymous):

282240?

OpenStudy (anonymous):

is this the correct answer?

OpenStudy (anonymous):

\[\frac{8!}{7!(8-7)!} = \frac{8!}{7!1!} = \frac{8!}{7!}\] With me so far?

OpenStudy (anonymous):

ohhhhh yess!!!! i put it all together! now i see!

OpenStudy (anonymous):

So what is the answer?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

you are so helpful!!!

OpenStudy (anonymous):

i have another question! can you help me?

OpenStudy (anonymous):

How many ways are there to seat 4 people in a 9-passenger automobile? Assume that some times there will be no one in the driver's seat.